The growth of Pseudomonas bacteria is modeled as a
Answer #1

a)

[A]= [A]0 e^-kt

Time t = 1.0 h = 60 min and [A]0= 1.0*10^3 cells /L and k = 0.035 min-1

[A]= [A]0 e^-kt

[A]= 1.0*10^3 cell / Le^-0.035 min-1 *60 min

[A]= 1.0*10^3 cell / Le^-2.1

[A]= 1.22*10^2 cell / L

(b)

ln[A]= -kt + ln[A]0

ln4.0*10^3 cells /L = ln 1.0*10^3 cells /L -kt

8.294 = 6.91 -0.035 min-1 t

0.035 min t =1.386

t=39,61 min